OpenStudy (anonymous):

find the y- intercept. x+4y=-4

OpenStudy (aroub):

4y=-x-4 y=-x/4 -1 so the y-intercept is -1

OpenStudy (anonymous):

so is it 0,-1

OpenStudy (aroub):

yes! :D

OpenStudy (anonymous):

thanks